# Studying stellar populations in Omega Centauri with phylogenetics

### v1

@@ Abstract
Methods. These include the Gaussian Mixture Model and Neighbor-Joining clustering algorithms applied to a set of chemical abundances of $\omega$ Cen  stellar members.
Results. Instead of using the classical approach in astronomy of grouping them into separate populations, we focused on how the stars are related to each other. In this way, we could identify stars that likely formed in globular clusters versus those originating from prolonged in-situ star formation and how these stars interconnect.
@@ 1 Introduction
Each chemical element is created inside stars, or during stellar explosions, and ejected into the interstellar medium on varying timescales and in different amounts (Kobayashi et al., 2020). These ejected elements can then become locked back into new stars, formed from this chemically-enriched interstellar gas. The chemical abundances of low-mass (long-lived) stars therefore serve as a fossil record of their birth environments, and, as such, are essential for studying galactic evolution (Freeman & Bland-Hawthorn, 2002; Tissera et al., 2012). Critically, this information traces all the key nucleosynthetic element-creation paths, namely $\alpha$-capture, iron-peak, and various neutron-capture modes. Thus, the more chemical elements we have available for analysis in each star in these fossil records, the better we can disentangle the evolutionary timescales and processes (Griffith et al., 2024). Today, there are many different chemical abundance measurements of thousands of individual stars in $\omega$ Cen published (Johnson & Pilachowski, 2010; Mészáros et al., 2021). From these datasets, many studies have attempted to use clustering methods to find groups in these stars and chemical elements to interpret the main characteristics (Mészáros et al., 2021; Alvarez Garay et al., 2024), yet no definitive conclusions regarding the history of $\omega$ Cen’s populations have been reached. One reason might be because these studies are rather focused to separate the stars into distinct groups, and not necessarily designed to study how these groups connect in an evolutionary and historical sense and are thus related to each other in an astrophysically viable evolutionary sequence.
@@
Biology has an extensive track record developing and implementing phylogenetic tools to study relationships among species and so reconstruct their histories. Phylogenetic trees study the accumulation of small modifications passed from one generation to the next to build a model of their relationships as they developed over time (Baum et al., 2005). The key assumption in the use of trees is that there is a level of heritability between generations. In biology that heritability is provided by DNA. Indeed, it is the fact that DNA is shared across all organisms that allows the reconstruction of relationships across all of life, and the expansion of genomics techniques has revolutionised phylogenetic methods (Yang, 2014). DNA carries information about shared ancestry, and large genomic datasets have been used to model and analyze trees in evolutionary biology.
@@
However, although DNA is the basis for most phylogenetic reconstructions in biology, as long as there is a mechanism of heritability then other systems can be used. All that is required is some means by which information (and traits) are transmitted from one generation to another. For example, languages are inherited (and subject to modification from generation to generation), and have been analysed phylogenetically (Bromham et al., 2022). Archaeologists increasingly use these techniques to reconstruct cultural history from non-biological materials (O’Brien et al., 2025). Phylogenetic methods have also been applied to nearby stars using chemical abundances as heritable traits (Jofré et al., 2017). Indeed, stars generate new chemical elements themselves, which are expelled into the interstellar medium upon their death. New stars form from this chemically enriched interstellar medium, inheriting its chemical pattern. This fullfills the heritability condition required for phylogenetic methods. Some low-mass stars survive for as long as the age of the Universe, and their chemical pattern serves as the fossil record and heritable marker of the evolving interstellar medium (Freeman & Bland-Hawthorn, 2002). It is thus possible to build on approaches from evolutionary biology, adapt tools to reconstruct the history of galaxies by using the chemical patterns of stars as the heritable marker, and illustrate the power and opportunity of such inter-disciplinary approach (Jofré et al., 2017; Jackson et al., 2021; de Brito Silva et al., 2024). The complex nature of $\omega$ Cen, and the wealth of available stellar abundance data, make it an ideal case for such research.
@@
In Sect. 2 we present the data used in this work, which we take from published catalogues. In Sect. 3 we explain the phylogenetic methods used to analyse $\omega$ Cen, and in Sect. 4 we provide an astrophysical interpretation of these results. We conclude our work in Sect. 5.
@@ 2 Data
In this work we use two published datasets of stars from $\omega$ Cen, both containing data of chemical abundances obtained from high resolution spectral analysis. The idea to use both datasets here is to further investigate the impact of our conclusions due to the choice of data. These datasets come from (Johnson & Pilachowski, 2010, hereafter Optical) and (Mészáros et al., 2021, hereafter Infrared) and we explain them with more detail below. Additionally, we checked that both the Optical and Infrared sets with detailed abundances are highly probable members of $\omega$ Cen, according to Vasiliev & Baumgardt (2021). This is based on the updated Gaia EDR3 kinematic properties of Milky Way clusters. Both datasets (IDs, brightness, and abundances) are listed in the tables of the supplementary material.
Figure 1: Color magnitude diagram of both optical (in blue) and IR (in red) selected samples. The stars used in this analysis are red giant branch stars.
@@ 2.1 Optical sample
Johnson & Pilachowski (2010) published stellar parameters and abundances for 855 red giants determined from spectra taken from Cerro Tololo Inter-American Observatory (CTIO) using the Blanco 4 m telescope equipped with the Hydra multifiber positioner and bench spectrograph. The spectra cover the wavelength range of $6135-6365$  and $\sim 6500-6800$, and have a resolving power of $R\sim 18,000$. The long exposure of the observations yielded high signal-to-noise spectra of about 200. The authors performed a spectroscopic analysis based on equivalent width methods to determine abundances of Fe, Na, Si, Ca, Sc, Ti, Ni and Eu. Abundances of O, Al, and La were determined using synthesis fitting. For Sc abundances of both ScI and ScII were provided, which we take the mean of. Among these 855 stars, 122 stars have abundances measured for all the elements. We consider this subsample for our analysis.
Photometry, coordinates, and membership probabilities for all stars in that study were taken from van Leeuwen et al. (2000). This is before the Gaia era, thus the proper motions were derived using photographic plates from the years 1931, 1937, 1978 and 1983. Because of this, obtaining updated photometry and astrometric information from Gaia for these stars was possible for only 710 stars, including all our 122 selected stars. In Fig. 1 we display with blue squares the colors and magnitudes of the stars from this sample. They follow the red giant branch, except a few outliers (see the study of multiple red giant branches in $\omega$ Cen by Pancino et al., 2000).
Figure 2: Abundance ratios against [Fe/H] of all elements in both optical (in blue) and IR (in red) samples. Some of the element abundances are measured in both samples, and some are measured in only one of the samples.
@@ 2.2 IR sample
This sample is taken from Mészáros et al. (2021), who analyzed stars observed with SDSS-IV/APOGEE-2 (Ahumada et al., 2020). It consists of 775 stars with APOGEE observations, which use an infrared spectrograph in the H band of a resolving power of $R\sim 22,000$ covering a range of 1.514-1.696 micrometers installed in the du Pont Telescope at Las Campanas Observatory. The data for $\omega$ Cen has a signal-to-noise of about 70. Using the code BACCHUS (Masseron et al., 2016), Mészáros et al. (2021) published abundances of Fe, C, N, O, Mg, Al, Si, K, Ca, and Ce. The subset which has measurements for all elements contains 137 stars and this is the set we use for our analysis. In this work we remove C and N abundances because in red giants they change as stars experience their dredge up (Salaris et al., 2015; Mészáros et al., 2021), making the photospheric abundances of these stars not a suitable heritable tracer.
Cluster membership for these stars is based on RV and proper motions from Gaia DR2 (Baumgardt & Hilker, 2018). The cross-match of this sample with Gaia DR3 yields 676 stars, with the 137 stars with detailed abundances included. Their position in the color-magnitude diagram can be seen in red circles in Fig. 1. As the Optical sample, they follow the red giant branch. There are few stars in the IR sample which are hotter and of lower luminosity compared to the Optical sample, but both datasets are otherwise comparable.
@@ 2.3 Chemical abundances
With the exception of a few stars from the Optical reaching values up to $\mathrm{[Fe/H]}=-0.3$ dex, both samples cover a comparable metallicity range between $-2$ and $-1$ dex. Only for O, Al, Si and Ca we have abundances for both samples. All these elements are produced in core-collapse Type II supernova (Nomoto et al., 2013), although Al is very dependent on the progenitor metallicity (Weinberg et al., 2019; Vasini et al., 2024) and has a different behaviour compared to the $\alpha-$capture element like O, Si, and Ca. Mg, which is another $\alpha-$capture element is only included in the IR sample, while Ti which shows the trends of an $\alpha-$capture element in the Milky Way, is only included as part of the optical sample. Ca, Mg, and Si, the other $\alpha-$capture elements of the sample, show a flatter trend as a function of metallicity.
@@
Al, Si and Ca agree well in both samples, showing similar trends and absolute values. This is not the case for O, which shows a positive trend with [Fe/H] and very high abundances in the Infrared sample, while the [O/Fe] trend with [Fe/H] decreases with [Fe/H] in the Optical sample. The high [O/Fe] abundances of some stars from the Infrared sample and its increasing trend with metallicity have also seen in the independent analysis of Marino et al. (2012), who analysed a sample of $\omega$ Cen stars observed with optical spectra. We recall that it is not expected that the plausible NSC nature of $\omega$ Cen has similar abundance patterns as the Milky Way or disrupted dwarf galaxies. Indeed, Romano et al. (2007) predicts high level of [O/Fe] abundance ratios of $\omega$ Cen, although with a decreasing trend with metallicity. Star formation bursts can cause an increase of [O/Fe] in dense systems, this was recently shown in the EDGE simulation of a NSC by Gray et al. (2024). We keep thus all [O/Fe] for our analysis.
@@
The Optical sample includes, in addition to Fe, the iron-peak abundance ratios Sc and Ni. Both Sc and Ni show flat trends and a small dispersion with [Fe/H], although Sc displays a small decrease over metallicity as it is being mostly produced by core-collapse supernovae (Kobayashi et al., 2020). Other elements produced in massive stars, which are only included in the Optical sample are Na and Eu. Both show trends with [Fe/H] and a high dispersion. Finally, both samples have one neutron--capture element produced by the $s-$process and expelled to the interstellar medium via asymptotic giant branch winds (see for instance Lugaro et al., 2003, 2012; Cseh et al., 2018; Magrini et al., 2018). The Optical sample includes La, and the Infrared sample includes Ce. Both elements show comparable trends, e.g. increasing with [Fe/H] from about -0.5 to 1 dex. This is expected (Bisterzo et al., 2014) because such elements produced in asymptotic giant branch stars are typically released after SNIa elements. Their trends indicates an extended star formation history.
@@ 2.4 Comparison of common stars between samples
Our cross-match with Gaia allows us to identify stars in common between both samples, finding 16 stars. Their IDs and Gaia magnitudes are listed in Table 1. In addition to the Source ID from Gaia, we include the APOGEE ID used by Mészáros et al. (2021), the LEID ID used in Johnson & Pilachowski (2010), and our own IDs which we call SIR for the Infrared sample, and SOP for the Optical sample.
The direct comparison for the abundances in common can be found in Fig. 3. The x-axis show the abundances as reported by Mészáros et al. (2021) and the y-axis show the abundances of the same stars as reported by Johnson & Pilachowski (2010) for [Fe/H], [O/Fe], [Al/Fe], [Si/Fe] and [Ca/Fe]. We overplot the one-to-one line in red and write in each panel the mean of the differences and the standard deviation of the differences. We note the small errors reported for Al. Indeed, for most stars, no Al errors are reported in Johnson & Pilachowski (2010). This is because in that work errors are the scatter among lines. Measuring Al from optical spectra is very challenging (see e.g. Buder et al. (2022)), the Al abundances here come mostly form one line which is why no errors were reported. While the agreement for Al, Si, and Ca are within the scatter, there is a systematic offset of 0.15 dex in metallicity between both samples, and a significant offset in [O/Fe] of 0.4 dex. The large difference in [O/Fe] between optical and infrared measurements in metal-poor stars has been already identified in the literature (Griffith et al., 2019). While applying a shift might bring into better agreement both samples in the [Fe/H] - [O/Fe] plane, the trends still would go on opposite directions. It is thus not possible to fully attribute the differences of oxygen between samples to a systematic effect, the difference might also be due to a selection effect of stars.
@@ 3 Phylogenetic analysis
Our goal is to use chemical abundances to study the assemblage of stars in $\omega$ Cen. This is analogous to biologists studying the genealogical history of a biological population, such as asking if a human population is descended from one or multiple sources through migrations and admixture (Hellenthal et al., 2014). This approach that biologists commonly use can be adapted to study the ancestry of stars. This process usually starts with a principal component analysis (PCA) to visualize the variability in the data, followed by a mixture model to detect distinct groups in the data, and finalised by a phylogenetic analysis to reconstruct when and where migration and admixture events happened (Nespolo et al., 2020; Price et al., 2006; Tian et al., 2008; Solovieff et al., 2010). In this study, we adapt this approach to help us understand the formation of $\omega$ Cen. We focus our analysis and discussion on the results obtained from Optical dataset, but the parallel results for the Infrared sample can be found in the appendixes.
@@
First, we apply PCA to our dataset. This is to ensure that there is chemical diversity in the dataset to detect distinct groups Buckley et al. (2024) and to understand which elements are responsible for this diversity. PCA has been done with chemical data of Milky Way stars (e.g. Ting et al., 2012; Andrews et al., 2012; Signor et al., 2024; Buckley et al., 2024; Griffith et al., 2024). In Fig. 4 we show the distribution of the 122 $\omega$ Cen stars in the space defined by the first three principle components (PCs), and the contribution of significant chemical abundance to each PC. The left-hand panel shows the first and second dimension, while the right-hand panel shows the first and the third dimension. A similar plot but for the Infrared sample can be found in Fig. 11. Each star is represented by a gray circle. The arrows of each abundance ratio are overplotted in the diagrams whose length and color correspond to the contribution in these dimensions while their direction of the vector correspond to where the stars have a higher or lower abundance ratio in each PCA dimension. These planes in a way summarize Fig. 5 of Buckley et al. (2024), who plotted the first and second dimension of the PCA color coded by different abundance ratios in 9 panels.
@@
The plane shows that an anticorrelation of aluminum and oxygen, as well as sodium and oxygen, have an important impact on the variance of the data. This can be seen through the [Na/Fe] and [Al/Fe] vectors pointing in the opposite direction to the [O/Fe] vectors. Indeed, the Na-O and Al-O anticorrelations are typical of GCs (Carretta et al., 2009), whose nature remains a mistery (Bastian & Lardo, 2018), and have been studied in $\omega$ Cen (Johnson & Pilachowski, 2010; Alvarez Garay et al., 2024). We see long arrows for those elements pointing towards opposite directions. The first PC captures 34.8% of the variability among stars in their chemical composition, which is mostly driven by the Na-O anticorrelation, with some influence from La, Si, Ca, Ti and Fe (see Sect. B.1). The second PC (Y axis of left hand panel), captures 17.9% of variance and separates the stars mostly by their $\alpha$-capture (O, Ca), iron-peak (Ni) elements and aluminum. The third dimension (Y axis of right-hand panel) contributes 14.9% to the variance and separates the stars in neutron-capture (Eu) abundance ratios as well as iron. For more in-depth analysis of the PCA, as well as the PCA using the Infrared sample, we refer the reader to the Appendix B.1. We conclude that we need to consider all available elements because they all contribute to one of the three most significant principal components.

### v2

@@ Abstract
Methods. These include the Gaussian mixture model and neighbor-joining clustering algorithms applied to a set of chemical abundances of $\omega$ Cen  stellar members. Instead of using the classical approach in astronomy of grouping them into separate populations, we focused on how the stars are related to each other.
Results. We could identify stars that likely formed in globular clusters versus those originating from prolonged in-situ star formation and how these stars interconnect.
@@ 1 Introduction
Chemical elements created inside stars, or during stellar explosions, and ejected into the interstellar medium on varying timescales and in different amounts (Kobayashi et al., 2020; Cowan et al., 2021). These ejected elements can then become locked back into new stars, formed from this chemically-enriched interstellar gas. The chemical abundances of low-mass (long-lived) stars therefore serve as a fossil record of their birth environments, and, as such, are essential for studying galactic evolution (Freeman & Bland-Hawthorn, 2002; Tissera et al., 2012). Critically, chemical abundances in stars trace all the key nucleosynthetic element-creation paths, namely $\alpha$-capture, iron-peak, and various neutron-capture modes. Thus, the more chemical elements we have available for analysis in each star in these fossil records, the better we can disentangle the evolutionary timescales and processes (Griffith et al., 2025). Today, there are many different chemical abundance measurements of thousands of individual stars in $\omega$ Cen published (Johnson & Pilachowski, 2010; Mészáros et al., 2021). From these datasets, many studies have attempted to use clustering methods to find groups in these stars and chemical elements to interpret the main characteristics (Mészáros et al., 2021; Alvarez Garay et al., 2024), yet no definitive conclusions regarding the history of $\omega$ Cen’s populations have been reached. One reason might be because these studies are rather focused to separate the stars into distinct groups, and not necessarily designed to study how these groups connect in an evolutionary and historical sense and are thus related to each other in an astrophysically viable evolutionary sequence.
@@
Biology has an extensive track record developing and implementing phylogenetic tools to study relationships among species and so reconstruct their histories. Phylogenetic trees study the accumulation of small modifications passed from one generation to the next to build a model of their relationships as they developed over time (Baum et al., 2005). The key assumption in the use of trees is that there is a level of heritability between generations. In biology that heritability is provided by DNA. Indeed, it is the fact that DNA is shared across all organisms that allows the reconstruction of relationships across all of life, and the expansion of genomics techniques has revolutionized phylogenetic methods (Yang, 2014). DNA carries information about shared ancestry, and large genomic datasets have been used to model and analyze trees in evolutionary biology.
@@
However, although DNA is the basis for most phylogenetic reconstructions in biology, as long as there is a mechanism of heritability then other systems can be used (O’Brien et al., 2025). All that is required is some means by which information (and traits) are transmitted from one generation to another. For example, languages are inherited (and subject to modification from generation to generation), and have been analyzed phylogenetically (Bromham et al., 2022). Archaeologists increasingly use these techniques to reconstruct cultural history from non-biological materials (O’Brien et al., 2025). Phylogenetic methods have also been applied to nearby stars using chemical abundances as heritable traits (Jofré et al., 2017). Indeed, stars generate new chemical elements themselves, which are expelled into the interstellar medium upon their death. New stars form from this chemically enriched interstellar medium, inheriting its chemical pattern. This fullfills the heritability condition required for phylogenetic methods. Some low-mass stars survive for as long as the age of the Universe, and their chemical pattern serves as the fossil record and heritable marker of the evolving interstellar medium (Freeman & Bland-Hawthorn, 2002). It is thus possible to build on approaches from evolutionary biology, adapt tools to reconstruct the history of galaxies by using the chemical patterns of stars as the heritable marker, and illustrate the power and opportunity of such inter-disciplinary approach (Jofré et al., 2017; Jackson et al., 2021; de Brito Silva et al., 2024). The complex nature of $\omega$ Cen, and the wealth of available stellar abundance data, make it an ideal case for such research.
@@
In Sect. 2 we present the data used in this work, which we take from published catalogs. In Sect. 3 we explain the phylogenetic methods used to analyse $\omega$ Cen, and in Sect. 4 we provide an astrophysical interpretation of these results. We conclude our work in Sect. 5.
@@ 2 Data
In this work we use two published datasets of stars from $\omega$ Cen, both containing data of chemical abundances obtained from high resolution spectral analysis. The idea to use both datasets here is to further investigate the impact of the choice of data on our conclusion. These datasets come from (Johnson & Pilachowski, 2010, hereafter Optical) and (Mészáros et al., 2021, hereafter Infrared) and we explain them in more detail below. Additionally, we checked that both the Optical and Infrared datasets with detailed abundances are highly probable members of $\omega$ Cen, according to Vasiliev & Baumgardt (2021). This is based on the updated Gaia EDR3 kinematic properties of Milky Way clusters. Both datasets (IDs, brightness, and abundances) are listed in the CDS tables.
Figure 1: Abundance ratios against [Fe/H] of all elements in both optical (in blue squares) and IR (in red circles) samples. Some of the element abundances are measured in both samples, and some are measured in only one of the samples. Shaded contours correspond to the Gaussian kernel density estimation of the datapoints.
@@ 2.1 Optical sample
Johnson & Pilachowski (2010) published stellar parameters and abundances for 855 red giants determined from spectra taken from Cerro Tololo Inter-American Observatory (CTIO) using the Blanco 4 m telescope equipped with the Hydra multifiber positioner and bench spectrograph. The spectra cover the wavelength range of $6135-6365$  and $\sim 6500-6800$, and have a resolving power of $R\sim 18,000$. The long exposure of the observations yielded high signal-to-noise spectra of about 200. The authors performed a spectroscopic analysis based on equivalent width methods to determine abundances of Fe, Na, Si, Ca, Sc, Ti, Ni, and Eu. Abundances of O, Al, and La were determined using synthesis fitting. For Sc abundances of both ScI and ScII were provided, which we take the mean of. Among these 855 stars, 122 stars have abundances measured for all the elements. We consider this subsample for our analysis.
Photometry, coordinates, and membership probabilities for all stars in that study were taken from van Leeuwen et al. (2000). This is before the Gaia era, thus the proper motions were derived using photographic plates from the years 1931, 1937, 1978, and 1983. Because of this, obtaining updated photometry and astrometric information from Gaia for these stars was possible for only 710 stars, including all our 122 selected stars. The stars are in the red giant branch, except a few outliers (see the study of multiple red giant branches in $\omega$ Cen by Pancino et al., 2000).
@@ 2.2 IR sample
This sample is taken from Mészáros et al. (2021), who analyzed stars observed with SDSS-IV/APOGEE-2 (Ahumada et al., 2020). It consists of 775 stars with APOGEE observations, which use an infrared spectrograph in the H band with a resolving power of $R\sim 22,000$ covering a range of 1.514-1.696 micrometers installed in the du Pont Telescope at Las Campanas Observatory. The data for $\omega$ Cen has a signal-to-noise of about 70. Using the code BACCHUS (Masseron et al., 2016), Mészáros et al. (2021) published abundances of Fe, C, N, O, Mg, Al, Si, K, Ca, and Ce. The subset which has measurements for all elements contains 137 stars and this is the set we use for our analysis. In this work we remove C and N abundances because in red giants they change as stars experience their dredge up (Salaris et al., 2015; Mészáros et al., 2021), making the photospheric abundances of these stars not a suitable heritable tracer.
Cluster membership for these stars is based on RV and proper motions from Gaia DR2 (Baumgardt & Hilker, 2018). The cross-match of this sample with Gaia DR3 yields 676 stars, with the 137 stars with detailed abundances included. As the Optical sample, they follow the red giant branch. There are a few stars in the Infrared sample which are hotter and of lower luminosity compared to the Optical sample, but both datasets are otherwise comparable.
@@ 2.3 Chemical abundances
With the exception of a few stars from the Optical reaching values up to $\mathrm{[Fe/H]}=-0.3$ dex, both samples cover a comparable metallicity range between $-2$ and $-1$ dex. We only have abundances of O, Al, Si, and Ca for both samples. All these elements are mainly produced in core-collapse Type II supernova (Nomoto et al., 2013), although Al is very dependent on the progenitor metallicity (Weinberg et al., 2019; Vasini et al., 2024) and has a different behaviour compared to the $\alpha-$capture elements like O, Si, and Ca. Mg, which is another $\alpha-$capture element is only included in the IR sample, while Ti, which shows the trends of an $\alpha-$capture element in the Milky Way, is only included as part of the optical sample. Ca, Mg, and Si, the other $\alpha-$capture elements of the sample, show a flatter trend as a function of metallicity. We remark that among these elements, O, Na, Mg, Al, and Si are those that show high levels of depletion or enhancement due to the internal stellar evolution processes of the CNO, NeNA, MgAl cycles that are suspected to produce the anticorrelations in GCs (Gratton et al., 2019).
@@
Al, Si, and Ca agree well in both samples, showing similar trends and absolute values. This is not the case for O, which shows a positive trend with [Fe/H] and very high abundances in the Infrared sample, while the [O/Fe] trend with [Fe/H] decreases with [Fe/H] in the Optical sample. This discrepancy is discussed further the next section. The high [O/Fe] abundances of some stars from the Infrared sample and its increasing trend with metallicity have also seen in the independent analysis of Marino et al. (2012), who analyzed a sample of $\omega$ Cen stars observed with optical spectra. We recall that it is not expected that the plausible NSC nature of $\omega$ Cen has similar abundance patterns as the Milky Way or disrupted dwarf galaxies. Indeed, Romano et al. (2007) predict high [O/Fe] abundance ratios in $\omega$ Cen, although with a decreasing trend with metallicity. Star formation bursts can cause an increase for [O/Fe] in dense systems, this was recently shown in the EDGE simulation of a NSC by Gray et al. (2025). We keep thus all [O/Fe] for our analysis.
@@
The Optical sample includes, in addition to Fe, the iron-peak abundance ratios Sc and Ni. Both Sc and Ni show flat trends and a small dispersion with [Fe/H], although Sc displays a small decrease over metallicity as it is being mostly produced by core-collapse supernovae (Kobayashi et al., 2020). Other elements produced in massive stars, which are only included in the Optical sample are Na and Eu. Both show trends with [Fe/H] and a high dispersion. Europium however is also produced by neutron star mergers (Cowan et al., 2021; Arcones & Thielemann, 2023), which are able to enrich entire dwarf galaxies with high levels of Eu in one episode of star formation. Finally, both samples have one neutron--capture element produced by the $s-$process and expelled to the interstellar medium via asymptotic giant branch (AGB) winds (see for instance Lugaro et al., 2003, 2012; Cseh et al., 2018; Magrini et al., 2018). The Optical sample includes La, and the Infrared sample includes Ce. Both elements show comparable trends, such as increasing with [Fe/H] from about -0.5 to 1 dex. This is expected (Bisterzo et al., 2014) because such elements produced in AGB stars are typically released after SNIa elements. Their trends indicates an extended star formation history.
@@ 2.4 Comparison of common stars between samples
Our cross-match with Gaia allows us to identify stars in common between both samples, finding 16 stars. Their IDs and Gaia magnitudes are listed in Table 1. In addition to the Source ID from Gaia, we include the APOGEE ID used by Mészáros et al. (2021), the LEID ID used in Johnson & Pilachowski (2010), and our own IDs, which we call SIR for the Infrared sample, and SOP for the Optical sample.
The direct comparison for the abundances in common can be found in Fig. 2. The x axis show the abundances as reported by Mészáros et al. (2021) and the y axis show the abundances of the same stars as reported by Johnson & Pilachowski (2010) for [Fe/H], [O/Fe], [Al/Fe], [Si/Fe], and [Ca/Fe]. We overplot the one-to-one line in red and write in each panel the mean of the differences and the standard deviation of the differences. We note the small errors reported for Al. Indeed, for most stars, no Al errors are reported in Johnson & Pilachowski (2010). This is because in that work errors are the scatter among lines. Measuring Al from optical spectra is very challenging (see for instance Buder et al. (2022)), the Al abundances here come mostly form one line, which is why no errors were reported. While the agreement for Al, Si, and Ca are within the scatter, there is a systematic offset of 0.15 dex in metallicity between both samples, and a significant offset in [O/Fe] of 0.4 dex. The large difference in [O/Fe] between optical and infrared measurements in metal-poor stars has been already identified in the literature (Griffith et al., 2019). While applying a shift might bring into better agreement both samples in the [Fe/H] - [O/Fe] plane, the trends still would go on opposite directions. It is thus not possible to fully attribute the differences in oxygen between samples to a systematic effect; the difference might also be due to a selection effect of stars.
@@ 3 Phylogenetic analysis
Our goal is to use chemical abundances to study the assembly history of stars in $\omega$ Cen. This is analogous to biologists studying the genealogical history of a biological population, such as asking if a human population is descended from one or multiple sources through migrations and admixture111 In population genetics, migration refers to the movement of genes from one population to another. This can occur when individuals migrate and introduce their genetic material into a new population. Admixture refers to the mixture of previously isolated populations. This can happen between different species (hybridization) or within the same species. (Hellenthal et al., 2014). This approach that biologists commonly use can be adapted to study the ancestry of stars. This process usually starts with a principal component analysis (PCA) to visualize the variability in the data, followed by a mixture model to detect distinct groups in the data, and finalized by a phylogenetic analysis to reconstruct when and where migration and admixture events happened (Nespolo et al., 2020; Price et al., 2006; Tian et al., 2008; Solovieff et al., 2010). In this study, we adapt this approach to help us understand the formation of $\omega$ Cen. We focus our analysis and discussion on the results obtained from the Optical dataset, but the parallel results for the Infrared sample can be found in the appendixes.
@@
First, we apply PCA to our dataset. This is to ensure that there is chemical diversity in the dataset to detect distinct groups Buckley et al. (2024) and to understand which elements are responsible for this diversity. PCA has been done with chemical data of Milky Way stars (Ting et al., 2012; Andrews et al., 2012; Signor et al., 2024; Buckley et al., 2024; Griffith et al., 2025). In Fig. 3 we show the distribution of 122 $\omega$ Cen stars in the space defined by the first three principle components (PCs), and the contribution of significant chemical abundances to each PC. The left-hand panel shows the first and second component, while the right-hand panel shows the first and the third component. A similar plot but for the Infrared sample can be found in Fig. 10. Each star is represented by a gray circle. The arrows of each abundance ratio are overplotted in the diagrams whose length and color correspond to the contribution in these dimensions while the direction of the vector corresponds to where the stars have a higher abundance ratio in each PC. These planes in a way summarize Fig. 5 of Buckley et al. (2024), who plotted the first and second component of the PCA color coded by different abundance ratios in nine panels.
@@
The plane shows that an anticorrelation of aluminum and oxygen, as well as sodium and oxygen, have an important impact on the variance of the data. This can be seen through the [Na/Fe] and [Al/Fe] vectors pointing in the opposite direction to the [O/Fe] vectors. Indeed, the Na-O and Al-O anticorrelations are typical of GCs (Carretta et al., 2009), whose nature remains unresolved (Bastian & Lardo, 2018; Gratton et al., 2019), and have been studied in $\omega$ Cen (Johnson & Pilachowski, 2010; Alvarez Garay et al., 2024). We see long arrows for those elements pointing to opposite directions. The first PC captures 34.8% of the variability among stars in their chemical composition, which is mostly driven by the Na-O anticorrelation, with some influence from La, Si, Ca, Ti, and Fe (see Sect. B.1). The second PC (Y axis of left hand panel), captures 17.9% of the variance and separates the stars mostly by their $\alpha$-capture (O, Ca), iron-peak (Ni) elements and aluminum. The third dimension (Y axis of right-hand panel) contributes 14.9% to the variance and separates the stars in neutron-capture (Eu) abundance ratios as well as iron. For more in-depth analysis of the PCA, as well as the PCA using the Infrared sample, we refer the reader to the Appendix B.1. When using the Infrared sample (see Fig. 10, the [O/Fe] and [Al/Fe] vectors point in the opposite directions to the Optical sample (reflecting the different trends with [Fe/H]), but the O-Al anticorrelation still holds. We conclude that we need to consider all available elements because they all contribute meaningfully to at least one of the three most significant principal components.